OpenStudy (anonymous):

Instead of finding 28% of a number Saloni incorrectly found 82% of the number. She got an answer that is 540 more than the answer that she is supposed to get. What is the answer she is supposed to get?

OpenStudy (anonymous):

28/100x-82/100x+x=540-x

OpenStudy (anonymous):

54/100x+x=540-x

OpenStudy (crashonce):

let x be the original number: we have 0.82x=0.28x+540 solve

OpenStudy (anonymous):

x=1000

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

nope

OpenStudy (dumbcow):

that is "x" not the answer

OpenStudy (dumbcow):

you found the number, they want to know what 28% of the number is

OpenStudy (anonymous):

1000-540=460

OpenStudy (anonymous):

1000*28/100=280

OpenStudy (dumbcow):

yes